OpenStudy (anonymous):

Find the coordinates of the vertex for the parabola defined by the given quadratic function. f(x)=5x^2-40x+22

OpenStudy (campbell_st):

well start by grouping things \[f(x) = (5x^2 + 40x) + 22\] remove the common factor from the brackets \[f(x) = 5(x^2 + 8x ) + 22\] you will need to complete the square for the expression in the brackets. you goal is to get it to the form \[y = a(x - h)^2 + k\] so what gets added into the brackets to make a perfect score...?

OpenStudy (campbell_st):

thats one option.... the alternate option is find the line of symmetry for a parabola \[y = ax^2 + bx + c\] the line of symmetry is \[x = \frac{-b}{2a}\] in your question b = 40 and a = 5 substitute them to find x... the vertex is on the line of symmetry... so substitute the x value you just found to get the y value of the vertex

OpenStudy (anonymous):

Ok I'll probably go with your first way of solving it!

OpenStudy (anonymous):

why did you only take 5 out instead of 5x?

OpenStudy (campbell_st):

well if you use the 2nd method \[x = \frac{-40}{2 \times 5}\] so x = - 4 this is the x- part of the vertex just substitute x = -4 to find the y part of the vertex..

OpenStudy (anonymous):

is this right? For the first method... 5(x+4)^2+38

OpenStudy (anonymous):

completing the square gave me (4+x)^2 than I added 16 to the 22

OpenStudy (campbell_st):

well almost (5(x + 4)^2 is great.... it expands to 5x^2 + 40x + 80... so to compensate you need to subtract 80 so you have y = 5(x + 4)^2 + 22 - 80 that keeps it in balance... so simplify the last bit to get it into vertex form.

OpenStudy (anonymous):

-4,-58?

OpenStudy (campbell_st):

yep... seems to make sense to me...

OpenStudy (anonymous):

-4 isn't right unfortunately

OpenStudy (anonymous):

I'm doing my homework and it says it's wrong.. but that -58 is right which doesn't make sense to me because I used -4 to get that

OpenStudy (anonymous):

Oh it's 4

OpenStudy (campbell_st):

no its -4.... unless you original equation is entered incorrectly... you can graph the curve using this site https://www.desmos.com/calculator and see the vertex is -4, -58

OpenStudy (campbell_st):

so there are a couple of options... the equation should be y = 5x^2 - 40x + 22 or... there is an error in entering the answer..... someone forgot to enter the negative when the answer was created.

OpenStudy (anonymous):

Maybe that was the case. So for this one, 2x^2-28x-144 I used your symmetry equation and got x=7

OpenStudy (anonymous):

then I plugged that in to get -242, but those are wrong. What did I do wrong?

OpenStudy (anonymous):

It is also asking for the x and y intercepts of that line

OpenStudy (campbell_st):

so x = 7 is correct so its 2(7)^2 - 28(7) - 144 = -242 thats correct as well well the y- intercept is y = -144 and the x intercepts are from factoring so y = 2(x^2 - 14x - 72) y = 2(x -18)(x + 4) so the x intercepts are 18, -4 are you going from a text or website...? because of you use something like wolframalpha http://www.wolframalpha.com/ it will get the vertex which is the same as the ones you have found

OpenStudy (campbell_st):

I'm confident, based on the 2 equations you posted, the vertex for each equation is correct. I'd question your teacher...

OpenStudy (anonymous):

how do i get the y intercepts?

OpenStudy (campbell_st):

y- intercepts are easy.... just let x = 0... and then you get y = -144

OpenStudy (anonymous):

The part I really struggle with is finding the domain and range for equations like this. How do I do that?

OpenStudy (campbell_st):

ok... so the domain.... what values of x can you put into the equation.. in both parabola... you can enter any value of x.... so the domain is all real x. the range is what are the output values ..... in both these parabolas they are concave up.... and the lowest value is the y part of the vertex.... so the range is y > vertex value the 2nd question should have a range y >= -242

OpenStudy (campbell_st):

I have to go to school... hope it helped... your answers are correct...

OpenStudy (anonymous):

thank you for your help! Take care!